# D-Band 2D MIMO FMCW Radar System Design for Indoor Wireless Sensing

### v3

@@ Abstract
In this article, we present an algorithm-hardware system design approach for the D-band 2D multi-input multi-output (MIMO) frequency-modulated continuous-wave (FMCW) radar in indoor wireless sensing. We propose a scaling approach for the SNR of the receiver (RX) and the output power of the transmitter (TX) with distance, which can significantly reduce power consumption of the radar array. 2D direction-of-arrival (DOA) estimation using the radar array in the presence of Gaussian noise is evaluated using the multiple signal classification (MUSIC) and the minimum-variance distortionless response (MVDR) algorithms, over 1–10 m indoor distance and 0–30 dB SNR of the RX. The MUSIC algorithm can provide higher accuracy and faster simulation for large arrays, for example more than 16 elements, while the MVDR algorithm outperforms for small arrays. Furthermore, we develop a hardware model for the 2D radar array to evaluate the power consumption of the TX, RX, and local oscillator (LO) distribution network. It is shown that the power consumption of the LO distribution network, which can dominate the power consumption of a large array operating in the D-band, can be reduced through a distribution strategy for the LO amplifiers within the LO network to compensate for passive losses. Performance of the D-band radar array is evaluated in the free-space and the through-wall indoor sensing scenarios. Specifically, a 64-element radar array using the MUSIC algorithm for 2D DOA estimation can achieve a root-mean-square error (RMSE) of $\rm{0.8^{\circ}}$/$\rm{0.3^{\circ}}$ at 10/20 dB SNR and 5 m distance. The required TX output power per radar element is estimated as $-$$-$16 dBm at 10/20 dB SNR in the free-space sensing. In the through-wall sensing, this increases to $-$10/0 dBm for clear glass, 3/13 dBm for drywall, and excessively high levels of 24/34 dBm for wood door, at 140 GHz.
Index Terms:
D-band, direction of arrival (DOA), internet of things (IoT), millimeter-wave, multi-input multi-output (MIMO), radar, receiver (RX), sixth generation (6G), transmistter (TX), wireless sensing.
Millimeter-Wave frequency bands above 100 GHz are promising for the next generation of wireless communication, sensing, and imaging systems. The extremely wide bandwidths available in these bands can enable data rates on the order of 100 Gbps for communication and millimeter-scale resolutions for sensing and imaging. The D-band, 110–170 GHz, is considered a potential candidate for sixth generation (6G) mobile communications and radar systems [1, 2, 3]. Wireless sensing with mm-wave radars is a key technology for emerging applications in integrated radar-communication systems, autonomous vehicles, gesture recognition for human-computer interaction, contactless health monitoring, and medical imaging.
Recently, several developments are presented towards the realization of wireless communications and sensing above 100 GHz. These developments, in the system and circuit levels, include experimental verification of the Friis free-space path loss (FSPL) model [2], characterization of the propagation channel [4], modeling and measurement of surface scattering [2], loss measurements of outdoor and indoor materials [2, 5], and implementation of integrated transceivers [3, 6, 7, 8, 9, 10, 11, 12, 13]. Silicon processes are preferred to III-V processes in most applications, despite the superior output power, efficiency, and noise performance of these processes (for example, GaN) [14], to achieve higher integration levels. State-of-the-art D-band silicon-based on-chip radars include 145 GHz frequency-modulated continuous-wave (FMCW) radar [6], 140 GHz FMCW radar with 25 GHz bandwidth [7], dual-mode W-band and D-band radar [9], 150 GHz joint radar-communication transceiver with 30 GHz bandwidth [10], and 155 GHz FMCW orthogonal frequency-division multiplexing (OFDM) radar with the chirp duration of 1–100 $\rm{\mu s}$ [11].
Multi-input multi-output (MIMO) radars can provide high angular resolutions, the spatial diversity, and improved effective SNR of the received signals [15, 16, 17, 18, 19]. The popular MIMO radar architecture based on 1D uniform linear array (ULA) can only detect azimuth angle which is insufficient for many emerging indoor wireless sensing applications. For example, in contactless heart monitoring, it is essential to focus the transmitted signal on the position of the human heart to receive accurate reflection signals. A 2D MIMO array should be used to detect azimuth and elevation angular information [17, 18, 19, 20], and combined with the range information perceived using FMCW signals, to realize 3D sensing/imaging systems.
The design and implementation of efficient D-band MIMO radars can be very challenging as a result of the multidisciplinary nature of the problem and many practical issues involved. However, the available approaches focus mainly on one aspect of the system design [3, 21, 22, 23, 24]. In the signal processing domain, many algorithms are developed to improve the resolution of directional-of-arrival (DOA) estimation, imaging, and perception. In the hardware level, there are several developments in antennas, RF circuits, transceivers, and digital signal processors, mainly focused to achieve higher performance, while the improvements and issues related to the signal processing which can have important effects on the hardware requirements are overlooked.
In this paper, we develop an algorithm-hardware system design approach for the D-band 2D MIMO FMCW radar in indoor wireless sensing. Main contributions of this work can be summarized as follows.
  1. 1.
    We develop an algorithm-hardware model to connect the accuracy of DOA estimation algorithms with the features of radar array hardware. The developed model is used to evaluate the performance of the radar array over 1–10 m distance in the free-space and through-wall indoor wireless sensing scenarios.
  2. 2.
    The direction finding accuracy of the 2D radar array is evaluated using two popular DOA estimation algorithms, multiple signal classification (MUSIC) [25] and minimum variance distortionless response (MVDR) [26], and it is shown that each of these algorithms can provide a superior accuracy and simulation speed under certain conditions.
  3. 3.
    We propose a scaling approach for the SNR of the radar receiver (RX) and the output power of the radar transmitter (TX) with distance. It is shown that this can significantly reduce the required TX output power and the power consumption of the radar array.
  4. 4.
    We develop a system-level model to evaluate the power consumption of the TX, RX, and LO distribution network of the radar array. The LO distribution network is shown to have a major contribution to the power consumption of the radar array, and a strategy is proposed to distribute the LO amplifiers across the network to compensate for losses with low power consumption overhead.
The paper is structured as follows. In Section I, the 2D MIMO FMCW radar model is presented. In Section II, the scaling approach for the SNR and output power is provided and the results of the 2D DOA estimation simulations are discussed. In Section III, a hardware model for the 2D radar array is developed to estimate the power consumption. In section IV, performance of the radar array is evaluated in the free-space and through-wall indoor wireless sensing scenarios. The investigation results and practical considerations are discussed in Section V. The conclusions are presented in Section VI.
I 2D MIMO FMCW Radar
I-A MIMO Radar Architecture
We consider a uniform array of $N_{x}\times N_{y}$ radar elements, as shown in Fig. 1. This array can be constructed as a virtual array comprising $N_{\rm{TX}}$ transmitters (TX) and $N_{\rm{RX}}$ receivers (RX). The number of virtual array elements is given by $N_{x}N_{y}=N_{\rm{TX}}N_{\rm{RX}}$. The distances are selected such that the inter-element spacing is $\lambda/2$. In the D-band, $\lambda/2\approx{\rm{1\,mm}}$, which allows large arrays to be integrated on a small physical area, for example, a 100-element array on the ${\rm{1\,cm\times 1\,cm}}$ area. This allows the radar array to be integrated into electronic devices (such as a modem) and provide wireless sensing capabilities.
I-B MIMO Waveforms
In the MIMO radar, orthogonal waveforms are transmitted to simplify the separation of received signals [27]. This can be achieved by applying a multiplexing scheme to modulate the chirp sequence signal $S_{\rm{TX}}(t)$ into the TX signals, $S_{\rm{TX,k}}(t)$ for $k=1,...,N_{\rm{TX}}$, as shown in Fig. 2.
Time-division multiplexing (TDM) is the simplest MIMO radar architecture, in which only one TX transmits at a time slot, that is $x_{k}(t)=p[(t-\Delta t_{k})/T]$, where $p(t)$ is the pulse function. However, this approach degrades the overall transmitted power, which is crucial for mm-wave bands. In the Doppler-division multiplexing (DDM) scheme, a phase code $x_{k}(t)=e^{j\Delta\Phi_{k}(t)}$ is applied with a specific pattern to realize different TX signals. In the orthogonal frequency-division multiplexing (OFDM) scheme, the TX signal is modulated by different carrier frequencies, which are usually shifted by an offset frequency $x_{k}(t)=e^{j[2\pi(k-1)\Delta f]t}$. In the DDM and OFDM MIMO architectures, all transmitters operate simultaneously and the total output power is enhanced. Therefore, we assume that the 2D MIMO array is realized using any of these multiplexing approaches, and any possible imperfections in the MIMO channels can be resolved through signal processing methods.
Figure 1: Architecture of 2D MIMO radar comprising $N_{x}\times N_{y}$ virtual arrays. For a uniform array operating in the D-band, $d_{x}=d_{y}=\lambda/2\approx{\rm{1\,mm}}$.
I-C Range Resolution
The FMCW radar transmits a frame of multiple chirp signals, correlates the received target reflection with the transmit signal, and detects the target range using the signal time of flight. The transmit signal for each chirp can be modeled by a complex chirp signal defined as
$S_{\rm{TX}}(t)=A_{\rm{TX}}e^{j(\omega_{0}t+\pi St^{2})}~{}~{}0\leq t\leq T_{c},$ (1)
where $\omega_{0}$ is the chirp start frequency, $S=B/T_{c}$ is the chirp slope, $B$ is the chirp bandwidth, and $T_{c}$ is the chirp duration. This signal is repeated in time intervals $(k-1)T_{c}\leq t\leq kT_{c}$ for $k=1,2,...,N_{c}$, where $N_{c}$ is the number of chirps in the frame (Fig. 2). The received signal can be derived as
$S_{\rm{RX}}(t)=A_{\rm{RX}}e^{j[\omega_{0}(t-\tau)+\pi S(t-\tau)^{2}]},$ (2)
where $\tau=2R/c$ is the radar signal time of flight and $R$ is the target distance. The received signal is mixed with a scaled version of the transmit signal, and is passed through a bandpass filter (BPF) to achieve the IF signal as
$S_{\rm{IF}}(t)=A_{\rm{IF}}e^{j(\omega_{0}\tau-\pi\tau^{2}S+2\pi\tau St)}.$ (3)
This indicates a sinusoidal signal with the frequency of
$f_{\rm{IF}}=\tau S=\frac{2R}{c}\frac{B}{T_{c}}.$ (4)
The radar range resolution is derived as the minimum distance between two close targets that their corresponding IF frequencies fall into two consecutive range bins
$\Delta R=\frac{c}{2B}.$ (5)
Therefore, the wide bandwidth (60 GHz) available in the D-band can offer a fine range resolution (2.5 mm).
Figure 2: Conceptual orthogonal waveform generation for MIMO FMCW radar. The chirp signal $S_{\rm{TX}}(t)$ is modulated by the sequence $x_{k}(t)$ to realize the transmit signals $S_{\rm TX,k}(t)$ for $k=1,...,N_{\rm{TX}}$.
I-D State-of-the-Art D-Band Radars
A performance summary of the state-of-the-art D-band on-chip FMCW radars is presented in Table I. A radar with 30 GHz bandwidth presented in [10] can provide a range resolution of 5.0 mm, which is promising for emerging sensing applications. Furthermore, ultrafast chirp signals, with a chirp sweep time of 1–100 $\rm{\mu s}$ and the slope of 1–8 GHz/$\rm{\mu s}$, can enable real-time high-resolution sensing scenarios. The TX output power is typically 10–13 dBm (10–20 mW) and the RX noise figure (NF) is 8–12 dB for the D-band on-chip radars. It should be noted that there is potential to further improve the radar performance in Table I. For example, the output power can be increased by combining output signals of multiple transistors.
TABLE I: State-of-the-art D-band on-chip FMCW radars.
@@ I-D State-of-the-Art D-Band Radars
(GHz) (GHz) $\rm{(\bm{\mu}\textbf{s})}$ (dBm) (dB)
[6] 145 10 5-30 $*\rm{{}^{*}}$ 8.0 28 nm CMOS
[7] 140 25 N/A 12.5 10.9 130 nm SiGe
[8] 122 20 N/A 13 N/A 130 nm SiGe
[9] 160 8 1000 13 N/A 130 nm SiGe
[10] 150 30 50 13 16 28 nm CMOS
[11] 155 8 1-100 9 7.5 22 nm FDSOI
[12] 122 10 1000 5 12.7 130 nm SiGe
[13] 145 18 80 $*\rm{{}^{*}}$ N/A 28 nm CMOS
$*\rm{{}^{*}}$ Effective Isotropic Radiated Power (EIRP)
I-E Angular Resolution
Figure 3: Magnitude of the array factor for different structures: (a) $8\times 8$ array, (b) $4\times 4$ array, (c) $2\times 2$ array.
@@ I-E Angular Resolution
where $\beta={2\pi}/{\lambda}$ is the wave number, $d_{x}$ and $d_{y}$ are the inter-element spacing in the $x$ and $y$ directions, and $(\theta,\phi)$ are the elevation and azimuth angles [28]. For the uniform array, $d_{x}=d_{y}=\lambda/2$. In Fig. 3, the magnitude of the array factor versus the azimuth and elevation angles is shown for different array sizes. A larger array provides higher directivity, narrower beamwidth, and lower side lobe level.
The angular resolution of the radar array can be derived as the separation between two consecutive nulls in the array factor $AF(\theta,\phi)$ versus the azimuth or elevation angle. The antenna elements have a limited beamwidth, which can be modeled as $|\phi|<\phi_{\max}$ and $|\theta|<\theta_{max}$. This leads to the angular resolutions in the azimuth and elevation as
$\Delta\phi\approx\frac{2}{N_{x}\cos{\phi_{max}}}$ (9)
$\Delta\theta\approx\frac{2}{N_{y}\cos{\theta_{max}}}.$ (10)
Therefore, the number of array elements should be selected based on the required angular resolutions and the beamwidth of the antennas. For example, assuming a beamwidth of ${}^{\circ}$ in both azimuth and elevation, the number of array elements to achieve an angular resolution of ${}^{\circ}$ should be $\rm{25\times 25}$.
In the presence of noise and interference signals, certain signal processing algorithms should be used to extract the signal features and estimate the DOA of the targets. The accuracy is dependent on the radar array size and structure, the SNR level, and the algorithm. Effective SNR of the MIMO radar under optimal signal combining is improved by a factor equal to the number of virtual array elements compared to SNR of each radar element

### v4

# Adaptive MIMO Radar Architecture for Energy-Efficient Wireless Sensing in the D-Band

@@ Abstract
I Introduction
Millimeter-Wave frequency bands above 100 GHz are promising for the next generation of wireless communication, sensing, and imaging systems. The extremely wide bandwidths available in these bands can enable data rates of multi 100 Gbps for communication and millimeter-scale resolutions for sensing and imaging. The D-band (110–170 GHz) is a candidate for 6G communications [1, 2, 3].
[…]
A multi-input multi-output (MIMO) radar is usually used in practical applications to provide sufficient angular resolutions, increase the spatial diversity, and improve effective SNR of the received signal [15, 16, 17, 18, 19, 20]. A major issue in conventional MIMO radars is the use of fixed hardware and algorithm for directional-of-arrival (DOA) estimation of the targets [3, 21, 22, 23, 24]. This leads to inefficient allocation of the radar resources in varying operational conditions.
We propose an adaptive MIMO radar architecture with the following key features: 1) the effective number of array elements is reconfigured based on the required accuracy of the DOA estimation to save the power consumption and reduce the computational complexity, 2) the output power of the TX and the required SNR of the RX are scaled with the distance to reduce the power consumption, and 3) the DOA estimation algorithm is selected based on the number of array elements to enhance accuracy and reduce the computational complexity. We develop a hardware model for the MIMO radar to estimate the power consumption of the TX, RX, and local oscillator (LO) distribution network. Performance of the adaptive MIMO radar is evaluated in the free-space and through-wall indoor wireless sensing scenarios, using two popular DOA estimation algorithms, multiple signal classification (MUSIC) [25] and minimum variance distortionless response (MVDR) [26].
The paper is structured as follows. In Section II, the adaptive MIMO radar architecture is presented. In Section III, the MIMO radar front-end is discussed and a circuit model is developed for the radar system to estimate the power consumption. In Section IV, the MIMO radar back-end including the DOA estimation algorithms is discussed. In section V, performance of the adaptive MIMO radar is evaluated for D-band wireless sensing. Finally, the concluding remarks are presented in Section VI.
II Adaptive MIMO Radar Architecture
II-A Adaptive Number of Array Elements
We consider a uniform 2D array of $N_{x}\times N_{y}$ radar elements, as shown in Fig. 1. This array can be constructed as a virtual array comprising $N_{\rm{TX}}$ transmitters and $N_{\rm{RX}}$ receivers. The number of virtual array elements is given by $N_{x}N_{y}=N_{\rm{TX}}N_{\rm{RX}}$. The distances are selected such that the inter-element spacing is $\lambda/2$. In the D-band, $\lambda/2\approx{\rm{1\,mm}}$, which allows large arrays to be integrated on a small physical area, for example, a 100-element array on a $\rm{1\,cm\times 1\,cm}$} area. This allows the MIMO radar to be integrated into electronic devices for wireless sensing capabilities.
The effective size of the radar array can be reconfigured by disabling certain TX and RX elements, as shown in Fig. 1. This decreases the angular resolution of the radar but also reduces the power consumption of RF circuits and the computational complexity of the digital signal processing (DSP) system. The key idea is to adaptively change the array size based on the required resolution in the operational conditions to increase energy efficiency of the sensing system.
II-B Adaptive Radar Resolution
A FMCW MIMO radar can detect 3D coordinates of the targets. The FMCW radar elements transmit a frame of multiple chirp signals, correlates the received target reflection with the transmit signal, and detects the target range using the signal time of flight. The angular information of the targets are extracted using the signals transmitted and received by the 2D MIMO array. The range resolution is dependent on the chirp signal bandwidth ($\Delta R=2c/B$), while the angular resolution can be adaptively controlled by the active number of array elements, transmitted power, and received SNR.
@@ II-B Adaptive Radar Resolution
where $\beta={2\pi}/{\lambda}$ is the wave number, $d_{x}$ and $d_{y}$ are the inter-element spacing in the $x$ and $y$ directions, and $(\theta,\phi)$ are the elevation and azimuth angles [28]. For a uniform array, $d_{x}=d_{y}=\lambda/2$. A larger array provides higher directivity, narrower beamwidth, and lower side lobe level. The angular resolution of the MIMO radar can be derived as the separation between two consecutive nulls in $AF(\theta,\phi)$ versus the azimuth or elevation angle. The antenna elements have a limited beamwidth, which can be modeled as $|\phi|<\phi_{m}$ and $|\theta|<\theta_{m}$. This leads to the angular resolutions as
$\Delta\phi\approx\frac{2}{N_{x}\cos{\phi_{m}}}$ (4)
$\Delta\theta\approx\frac{2}{N_{y}\cos{\theta_{m}}}.$ (5)
The number of array elements should be selected based on the required angular resolutions and the beamwidth of the antennas. In the adaptive MIMO radar with varying number of array elements based on the operating conditions, this helps to reduce the power consumption and the computational complexity.
In the presence of noise and interference signals, certain signal processing algorithms such as MUSIC should be used to extract the signal features and estimate the DOA. Effective SNR of the MIMO radar under optimal signal combining is improved by a factor equal to the number of virtual array elements compared to SNR of each radar element